# From Instruction to Event: Sound-Triggered Mobile Manipulation

### v1

@@ 3.2 Skill
Table 1: (a) Comparison of different simulators. Our Habitat-Echo supports sound rendering and physical interaction at the same time. (b) Distribution of audio assets across different categories in Habitat-Echo. We ensure a diverse split of sound instances for training and test sets.
@@
Split Alarm Furby Phone Sink Doorbell Total
Train 11 30 19 6 11 77
Test 5 14 9 3 5 36
Total 16 44 28 9 16 113
4 Simulator and Benchmark
4.1 Simulator: Habitat-Echo
Existing simulators (Kolve et al., 2017; Makoviychuk et al., 2021; Szot et al., 2021) prioritize physical interaction with visual observations, while others (Chen et al., 2020, 2022) incorporate audio-visual rendering specifically for navigation tasks (see Table 1(a)). There remains an absence of simulators that integrate sound rendering with physical interaction. We address this limitation by introducing Habitat-Echo, an extension of Habitat (Szot et al., 2021) that enables sound rendering alongside physical interaction.
@@ 4.1 Simulator: Habitat-Echo
To enable high-fidelity acoustic simulation within the environment, we leverage precomputed Room Impulse Responses (RIRs) in SoundSpaces (Chen et al., 2020). An RIR serves as an accurate transfer function characterizing the multipath propagation between a source and a receiver (Kuttruff, 2016). RIRs provide high-fidelity acoustic features by incorporating two critical physical factors: (1) Geometry-aware propagation: Generated via bidirectional path tracing, RIRs implicitly capture intricate wave phenomena, including direct paths, specular reflections, and reverberation. In this way, rendered sounds are consistent with the room’s 3D geometry. (2) Material-dependent scattering: The RIR computation accounts for acoustic material properties by mapping semantic annotations to frequency-dependent absorption and transmission coefficients. This enables the simulator to reflect distinct acoustic interactions with diverse surfaces. With the help of RIRs, Habitat-Echo synthesizes the auditory observation by performing time-domain convolution between the source waveform and the corresponding RIR on the fly. Furthermore, exploiting the property of the acoustic field, our simulator supports simultaneous multi-source rendering by superimposing the convolved signals from distinct sources, which is essential for the Bi-Sonic Manipulation task. The final output is rendered as binaural audio, providing the agent with realistic spatial cues analogous to human auditory perception.
4.2 Benchmark
@@
Sounding Objects and Audio Assets. In addition to standard YCB objects (Calli et al., 2015), we introduce five more categories of interactive sounding objects. For SonicStow, we introduce graspable rigid objects, including phone, alarm, and furby. These objects can be manipulated via Pick and Place skills. For SonicInteract, we incorporate articulated objects, including the sink and door. Unlike rigid items, these objects possess kinematic joints and are constrained to move along fixed functional axes. For the sink, we incorporate a dynamic water flow that is physically coupled to the faucet handle, moving synchronously with the handle’s rotation. To foster robust acoustic generalization, we also collect a diverse library of real-world sounds, implementing a strict one-to-many mapping where each object category is associated with multiple unique sound instances (see Table 1(b)). This design forces the agent to learn semantically meaningful audio-visual associations rather than overfitting to specific frequency patterns or waveforms, thereby ensuring high acoustic diversity of the benchmark.
@@
Episode Generation. We employ a procedural randomization process to generate episodes for training and testing. In each episode, the sound for the target object is determined by hierarchically sampling a sound category and then a specific audio instance. Visual complexity is introduced by instantiating random silent objects as distractors alongside the target object. The agent is initialized at a random position, and target objects (excluding two doorbells in the scene) are also placed at random positions within the scene. We also validate the solvability of every episode by ensuring a traversable path connects the start and goal, and we strictly verify that interactive objects have accessible frontal regions for valid manipulation, especially for sinks and doors. An episode is considered successful if and only if the required skills are successfully executed in sequence.
5 Baseline
@@
Figure 3: Overview of the proposed baseline. The sound-triggered task planner processes initial audio-visual observations to reason and generate a high-level skill chain from the skill library. Guided by this chain, specialized policy models are sequentially activated to generate low-level actions and interact with the environment.
@@
5.1 Sound-triggered Task Planner
@@
5.2 Policy Model
@@
Policy models serve as low-level executors, translating the skill chain from the task planner into concrete actions and sequentially interacting with the environment. Policy models in the proposed baseline share a similar architecture comprising modality encoders, a temporal module, and actor-critic heads. The temporal module and actor-critic modules maintain a consistent structure following prior works (Chen et al., 2020; Gu et al., 2023), and the modality encoders are tailored to the specific observation space of different skills. For the Navigate skill, the observation space includes audio and head RGB, and the modality encoders are designed based on task complexity. In SonicStow and SonicInteract, the audio observations are converted into spectrograms and processed by a vanilla CNN encoder (Chen et al., 2020). In Bi-Sonic Manipulation, audio observations are transformed into Mel-spectrograms, and the encoder is LanguageBind (Zhu et al., 2024). An additional text encoder is introduced to specify the target sound source. For manipulation skills, including Pick, Place, Open Door, and Close Sink, the observation spaces consist of head and arm depth, which are processed by a standard CNN encoder (Chen et al., 2020).
Table 2: Results on the Habitat-Echo benchmark. We report the success rate ($\%$) of task planning, individual skills, and overall performance across different tasks. The higher the better. (a) Evaluation of SonicStow. (b) Evaluation of SonicInteract. (c) Evaluation of Bi-Sonic Manipulation. Success rates are reported in the format “First Sound Source / Second Sound Source”.
Baseline Task Planning Navigate Pick Place Overall
Individual N/A 57.66 88.73 70.27 N/A
Oracle 100.0 57.66 86.72 81.98 40.99
Qwen2.5-Omni-3B 12.16 55.56 93.33 71.43 4.50
Qwen3-Omni-30B 68.92 55.56 87.06 72.97 24.33
Qwen2.5-Omni-7B 78.38 51.72 85.56 79.22 27.48
Method Task Planning Navigate Open Door Close Sink Overall
Individual N/A 45.50 62.16 77.67 N/A
Oracle 100.00 45.50 50.00 72.63 27.90
Qwen2.5-Omni-3B 44.14 3.00 50.00 0.00 0.33
Qwen3-Omni-30B 59.01 44.70 33.33 73.20 14.05
Qwen2.5-Omni-7B 71.17 44.63 50.00 74.16 19.72
Method Task Planning Navigate Pick Place Open Door Close Sink Overall
Individual N/A 9.58 / 56.61 85.16 / 17.39 32.48 / 6.00 67.12 / 58.66 42.87 / 38.00 N/A
Oracle 100.00 9.58 / 46.67 92.31 / 0.00 25.00 / 0.00 72.73 / 0.00 40.00 / 100.00 4.23 / 0.56
Qwen2.5-Omni-3B 14.08 8.00 / 0.00 100.00 / 0.00 0.00 / 0.00 0.00 / 0.00 100.00 / 0.00 0.56 / 0.00
Qwen2.5-Omni-7B 17.18 14.75 / 100.00 100.00 / 0.00 0.00 / 0.00 100.00 / 0.00 0.00 / 100.00 0.56 / 0.56
Qwen3-Omni-30B 51.83 11.96 / 0.00 100.00 / 0.00 0.00 / 0.00 66.67 / 0.00 37.50 / 0.00 2.53 / 0.00
Figure 4: Qualitative Visualization of Task Execution. We present the execution trajectories for (a) SonicStow, (b) SonicInteract, and (c) Bi-Sonic Manipulation. For each task, the left panel illustrates the top-down view of the environment, highlighting the agent’s trajectories and the location of sound sources. The right panel displays the corresponding sequence of third-person keyframes during the execution. Note that Bi-Sonic Manipulation involves navigating to and interacting with two distinct sound sources sequentially.
6 Experiment
6.1 Implementation
@@
Implementation Detail. We optimize all policy models using the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) on a single NVIDIA RTX 4090 GPU. The training process spans $5\times 10^{5}$ environment steps. Structurally, the visual and audio encoders remain frozen to ensure efficiency, and the Actor and Critic networks are Multi-Layer Perceptrons (MLPs) with a hidden dimension of 512. We utilize the Adam optimizer (Kingma, 2014) with an initial learning rate of $3\times 10^{-4}$ enabled with linear decay, a momentum of $1\times 10^{-5}$, and a maximum gradient norm of 0.5. Regarding PPO hyperparameters, we set the discount factor of 0.99, GAE parameter of 0.95, and clipping range of 0.2. The update phase involves 2 epochs with 2 mini-batches, using a rollout length of 128 steps. The value loss coefficient is set to 0.5, and the entropy coefficient is 0.1. For each task, the dataset is split into 660 instances for training. The testing dataset comprises 222 instances for both SonicStow and SonicInteract, and 355 instances for Bi-Sonic Manipulation.
@@
Baseline. We train low-level policy models individually and then compose them with varying task planners, implementing three types of baselines. First, the Individual baseline evaluates each skill in isolation. This baseline bypasses the task planner to measure each skill without the influence of cascading errors. Second, the Oracle baseline employs a privileged task planner that provides the ground-truth skill chains, serving as a performance upper bound for the task planning. Finally, we integrate mainstream Omni-LLMs as task planners to evaluate audio-visual reasoning capabilities, including Qwen2.5-Omni (3B, 7B) (Xu et al., 2025) and Qwen3-Omni-30B-Instruct (Yang et al., 2025). These task planners are all training-free. We refer to these baselines directly by their own names.
6.2 Result
Quantitative Result of SonicStow. We present the evaluation results on the SonicStow task in Table 2(a). The Oracle baseline, which utilizes ground-truth planning, establishes the performance upper bound with an overall success rate of 40.99%. The Individual baseline reflects the standalone success rate of each policy model (e.g., 88.73% for Pick), decoupling low-level execution and high-level planning. Regarding task planning, Qwen2.5-Omni-7B achieves the highest accuracy of 78.38%, surpassing both the smaller Qwen2.5-Omni-3B and the larger Qwen3-Omni-30B. Driven by this, Qwen2.5-Omni-7B generates the highest number of correct skill chains, thereby attaining the best overall execution performance of 27.48% among all task planners.
@@ 6.2 Result
Quantitative Result of SonicInteract. The evaluation results on the SonicInteract are shown in Table 2(b). With ground-truth task planning, the Oracle baseline achieves a 27.90% overall success rate, which is the theoretical upper bound of the policy models. Among the evaluated task planners, Qwen2.5-Omni-7B also emerges as the most effective one. It attains a planning rate of 71.17%, outperforming both Qwen2.5-Omni-3B and Qwen3-Omni-30B. The success rate of planning directly translates to execution success, and Qwen2.5-Omni-7B achieves the top overall performance of 19.72%. Conversely, we observe unsatisfactory cases in Qwen2.5-Omni-3B during the Close Sink subtask. Instead of responding to the sound of running water with the appropriate skill, Qwen2.5-Omni-3B consistently hallucinates a Pick and Place sequence, resulting in a near-zero success rate for this specific task.
@@
Quantitative Result of Bi-Sonic Manipulation. Table 2(c) presents the results for the Bi-Sonic Manipulation task. In this dual-source environment, Qwen3-Omni-30B achieves the highest task planning accuracy, showing superior robustness against acoustic distraction. In contrast, Qwen2.5-Omni-7B fails to maintain its dominance in the task planning, receiving performance comparable to Qwen2.5-Omni-3B. Regarding the overall success rate, the Oracle and Qwen3-Omni-30B secure the top two results for the first sound source interaction. For the secondary source interaction, only the Oracle and Qwen2.5-Omni-7B manage to achieve 2 successful episodes. This near-zero success rate highlights the significant challenges posed by the Bi-Sonic Manipulation task (see Section. 3.1).
Table 3: Ablation Study on Input Modalities for Different Policy Models. We investigate optimal input combinations across three representative policy models for different skills, including Navigate, Pick, and Open Door. SR denotes success rate. (a) Policy model for Navigate. We evaluate different modality encoders including DINOv2 (Oquab et al., 2024) and CNN encoder (Chen et al., 2020) for visual inputs. Results indicate that combining Head RGB and Head Depth yields the highest success rate. (b) Policy model for Pick. Using a same encoder, the combination of Head Depth and Arm Depth outperforms single-view inputs. (c) Policy model for Open Door. Similarly, integrating Head Depth and Arm Depth is also proven effective for achieving optimal performance in the Open Door skill.
Model Input Modality Modality Encoder SR ($\uparrow$)
Head RGB Head Depth DINOv2 CNN
I 57.66
II 58.01
III 51.80
IV 61.71
Model Input Modality SR ($\uparrow$)
Head Depth Arm Depth
I 57.20
II 66.66
III 88.73
Model Input Modality SR ($\uparrow$)
Head Depth Arm Depth
I 46.39
II 47.29
III 62.16
@@
Qualitative Result. Figure 4 presents qualitative visualizations of the agent’s execution across the three proposed tasks. In each subfigure, the left panel depicts the top-down trajectory and sound source locations, while the right panel displays sequential third-person keyframes of the interaction. Figure 4 (a) shows the SonicStow task. Initially, the target object (an alarm) is located outside the agent’s field of view. Relying on sound signals for localization, the agent successfully navigates to the target and executes the Pick and Place skills. Figure 4 (b) illustrates the SonicInteract task. The agent accurately navigates to the sound-emitting door and performs the Open Door skill. Figure 4 (c) showcases the Bi-Sonic Manipulation task, which involves two simultaneous sound sources: a primary source (yellow) and a secondary source (green). Guided by the task planner, the agent prioritizes the primary source despite the acoustic interference. It first navigates to and opens the door. Once the primary source is interacted, the agent isolates the remaining secondary signal, navigates to the sink, and executes the Close Sink skill. More visualizations can be seen in supp.
6.3 Ablation Study and Further Discussion

### v2

@@ 4.1. Habitat-Echo Simulation Platform
Table 1. Comparison of different simulators. Our Habitat-Echo supports sound rendering and physical interaction at the same time.
@@
As shown in  Table 1, existing simulators either prioritize visual-physical manipulation  (Kolve et al., 2017; Makoviychuk et al., 2021; Szot et al., 2021) prioritize physical interaction with visual observations, or strictly focus on audio-visual navigation (Chen et al., 2020, 2022), failing to support the audio-visual-physical closed-loop interaction. To bridge this gap, we develop Habitat-Echo.
Environment and Asset Composition. Habitat-Echo is built upon interactive indoor 3D environments (Szot et al., 2021) and extends them with a sound-emitting interactive asset library tailored for STMM. This library comprehensively covers the predefined interaction types, including graspable sounding rigid bodies (e.g., phones, alarms) and articulated or functional objects (e.g., doors, sinks with a faucet on), providing a robust asset foundation for both object relocation and state transition tasks. These assets are explicitly designed to act as both acoustic event sources and manipulation targets, thereby closing the loop between perception and action.
@@
Audio-Physical Simulation. To enable high-fidelity acoustic rendering, we leverage precomputed Room Impulse Responses (RIRs) in SoundSpaces (Chen et al., 2020). An RIR serves as an accurate transfer function characterizing the multipath propagation between a source and a receiver (Kuttruff, 2016). RIRs provide high-fidelity acoustic features by incorporating two critical physical factors: (1) Geometry-aware propagation: Generated via bidirectional path tracing, RIRs implicitly capture intricate wave phenomena, including direct paths, specular reflections, and reverberation. In this way, rendered sounds are consistent with the room’s 3D geometry. (2) Material-dependent scattering: The RIR computation accounts for acoustic material properties by mapping semantic annotations to frequency-dependent absorption and transmission coefficients. This enables the simulator to reflect distinct acoustic interactions with diverse surfaces. With the help of RIRs, Habitat-Echo synthesizes the auditory observation by performing time-domain convolution between the source waveform and the corresponding RIR on the fly. Furthermore, exploiting the property of the acoustic field, our simulator supports simultaneous multi-source rendering by superimposing the convolved signals from distinct sources, which is essential for the multi-event setting. The final output is rendered as binaural audio, providing the agent with realistic spatial cues analogous to human auditory perception.
Skill Interface. To support the skill space $\mathcal{K}$ defined in  Section. 3.3, Habitat-Echo provides standardized execution interfaces and underlying observation/action space support for each skill primitive (e.g., supporting Navigate termination based on audio-visual features, or OpenDoor operations bounded by joint constraints). The platform allows for the independent initialization, termination, and success detection of each skill, thereby enabling the sequential scheduling and closed-loop evaluation of entire skill chains.
4.2. STMM-1.9K Benchmark
Figure 3. (a) Distribution of audio assets. The semantic are shown in different colors, ranging from Electronic Doorbell Alert to Ambient Notifications. (b) Episode statistics. Train and test splits. O.R. denotes Object Relocation, and S.T. denotes State Transition.
@@
Dataset Construction. In addition to YCB objects (Calli et al., 2015), we introduce more categories of interactive sounding objects. For object relocation, we introduce graspable rigid objects, including phone, alarm, and furby. For state transition, we incorporate articulated objects, including the sink and door. Unlike rigid items, these objects possess kinematic joints and are constrained to move along fixed functional axes. For the sink, we incorporate a dynamic water flow that is physically coupled to the faucet handle, moving synchronously with the handle’s rotation. To foster robust acoustic generalization, we also collect a diverse library of real-world sounds, implementing a strict one-to-many mapping where each object category is associated with multiple unique semantics (see Figure 3(a)). This design forces the agent to learn semantically meaningful audio-visual associations rather than overfitting to specific frequency patterns or waveforms, ensuring high acoustic diversity of the benchmark.
@@
Episode Generation. We employ a procedural randomization process to generate episodes for training and testing. In each episode, the sound for the target object is determined by hierarchically sampling a sound category and then a specific audio instance. Visual complexity is introduced by instantiating random silent objects as distractors alongside the target object. The agent is initialized at a random position, and target objects (excluding two doorbells in the scene) are also placed at random positions within the scene. We also validate the solvability of every episode by ensuring a traversable path connects the start and goal, and we strictly verify that interactive objects have accessible frontal regions for valid manipulation, especially for sinks and doors. In the single-event setting, the dataset provides 660 training and 222 testing episodes for both object relocation and state transition. The multi-event setting contains 660 training and 355 testing episodes (See Figure 3(b)). An episode is considered successful strictly when task planning succeeds, and all prerequisite skills are completed in the correct order. Consequently, performance in the proposed setting is measured by success rates of task planning and skill completion.
@@
Figure 4. Overview of the proposed baseline. The sound-triggered task planner processes the initial observation to reason and generate a high-level skill chain from the skill library (right). Guided by this chain, specialized policy models are sequentially activated to generate low-level actions and interact with Habitat-Echo.
5. Hierarchical Baseline for STMM
@@
5.1. Sound-triggered Task Planner
@@
5.2. Skill-Specific Policy Model
@@
Policy models serve as low-level executors, translating the skill chain from the task planner into concrete actions and sequentially interacting with the environment. Policy models in the proposed baseline share a similar architecture comprising modality encoders, a temporal module, and actor-critic heads. The temporal module and actor-critic modules maintain a consistent structure following prior works (Chen et al., 2020; Gu et al., 2023), and the modality encoders are tailored to the specific observation space of different skills. For the Navigate skill, the observation space includes audio and head RGB, and the modality encoders (Chen et al., 2020; Oquab et al., 2024) are designed based on task complexity. In object relocation and state transition, the audio observations are converted into binaural spectrograms and processed by a vanilla CNN encoder (Chen et al., 2020). In the multi-event setting, audio observations are transformed into Mel-spectrograms, and the encoder is LanguageBind (Zhu et al., 2024). An additional text encoder is introduced to specify the target sound source. For manipulation skills, including Pick, Place, Open Door, and Close Sink, the observation spaces consist of head and arm depth, which are processed by a standard CNN encoder (Chen et al., 2020).
6. Experiment
6.1. Implementation
@@
Implementation Detail. We optimize all policy models using the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) on a single NVIDIA RTX 4090 GPU. The training process spans $5\times 10^{5}$ environment steps. Structurally, the visual and audio encoders remain frozen to ensure efficiency, and the Actor and Critic networks are Multi-Layer Perceptrons (MLPs) with a hidden dimension of 512. We utilize the Adam optimizer (Kingma, 2014) with an initial learning rate of $3\times 10^{-4}$ enabled with linear decay, a momentum of $1\times 10^{-5}$, and a maximum gradient norm of 0.5. Regarding PPO hyperparameters, we set the discount factor of 0.99, GAE parameter of 0.95, and clipping range of 0.2. The update phase involves 2 epochs with 2 mini-batches, using a rollout length of 128 steps. The value loss coefficient is set to 0.5, and the entropy coefficient is 0.1. For each task, the dataset is split into 660 instances for training. The testing dataset comprises 222 instances for both Object Relocation and State Transition, and 355 instances for Multi-event Setting.
@@
Baseline. We train low-level policy models individually and then compose them with varying task planners, implementing two types of baselines. First, the Oracle baseline employs a privileged task planner that provides the ground-truth skill chains, serving as a performance upper bound. Second, we integrate mainstream Omni-MLLMs as task planners to evaluate audio-visual reasoning capabilities, including Qwen2.5-Omni (3B, 7B) (Xu et al., 2025) and Qwen3-Omni-30B-Instruct (Yang et al., 2025). These task planners are all training-free. We refer to these baselines directly by their own names.
Method Single-event setting Multi-event Setting
State Transition Object Relocation
Planning ($\uparrow$) Completion ($\uparrow$) Planning ($\uparrow$) Completion ($\uparrow$) Planning ($\uparrow$) Completion ($\uparrow$)
Oracle 100.00 27.90 100.00 40.99 100.00 0.56
Oracle* 100.00 81.08 100.00 70.72 100.00 37.18
Qwen2.5-Omni-3B 44.14 0.33 12.16 4.50 14.08 0.00
Qwen3-Omni-30B 59.01 14.05 68.92 24.33 51.83 0.00
Qwen2.5-Omni-7B 71.17 19.72 78.38 27.48 17.18 0.56
Qwen2.5-Omni-3B* 44.14 29.27 12.16 8.55 14.08 3.09
Qwen3-Omni-30B* 59.01 53.15 68.92 48.19 51.83 19.71
Qwen2.5-Omni-7B* 71.17 59.00 78.38 53.60 17.18 4.78
Figure 5. Results on STMM-1.9k. (a) Success rates of task planning and skill completion. Qwen2.5-Omni-7B and Qwen3-Omni-30B receive the best performance on single-event and multi-event settings, respectively. (b, c) Progressive success rate of state transition completion. (d) Progressive success rate of object relocation completion. (e) Progressive success rate of skill completion in the multi-event setting. denotes the target position that is provided for the navigate policy model.
Figure 6. Qualitative Visualizations of Task Execution. We present the execution trajectories for (a) Object Relocation, (b) State Transition, and (c) Multi-event Setting. For each task, left panel illustrates the top-down view, highlighting the agent’s trajectories and the location of sound sources. The right panel displays the sequence of third-person keyframes. Note that multi-event setting involves navigating to and interacting with two distinct sound sources sequentially.
6.2. Result
@@
Quantitative Result of State Transition in Single-event setting. The evaluation results on the State Transition are shown in Figure 5 (a-c). Figure 5(a) shows the numerical performance of task planning and skill completion. Figure 5(b) and Figure 5(c) visualize the progressive success rate of State Transition. denotes the target position that is provided for the navigate policy model. With ground-truth task planning, the Oracle baseline achieves an 81.08% overall success rate, which is the theoretical upper bound of the policy models. Among the evaluated task planners, Qwen2.5-Omni-7B also emerges as the most effective one. It attains a planning rate of 71.17%, outperforming both Qwen2.5-Omni-3B and Qwen3-Omni-30B. The success rate of planning directly translates to execution success, and Qwen2.5-Omni-7B and Qwen2.5-Omni-7B achieve the top overall performance of 59.00% and 19.72%, respectively. Conversely, we observe unsatisfactory cases in Qwen2.5-Omni-3B during the Close Sink subtask. Instead of responding to the sound of running water with the appropriate skill, Qwen2.5-Omni-3B consistently hallucinates a Pick and Place sequence, resulting in a near-zero success rate for this specific setting.
Quantitative Result of Object Relocation in Single-event setting. We present the evaluation results on the Object Relocation in Figure 5 (a, d). The Oracle baseline, which utilizes ground-truth planning and target position, establishes the performance upper bound with an overall success rate of 70.72%. Regarding task planning, Qwen2.5-Omni-7B achieves the highest accuracy of 78.38%, surpassing both the smaller Qwen2.5-Omni-3B and the larger Qwen3-Omni-30B. Driven by this, Qwen2.5-Omni-7B generates the highest number of correct skill chains, thereby attaining the best skill completion performance of 53.60% among all.
@@
Quantitative Result of Multi-event Setting. Figure 5 (a, e) presents the results for the multi-event setting. In this multi-event environment, Qwen3-Omni-30B achieves the highest task planning accuracy, showing superior robustness against acoustic distraction. In contrast, Qwen2.5-Omni-7B fails to maintain its dominance in the task planning, receiving performance comparable to Qwen2.5-Omni-3B. For skill completion, Oracle sets an upper bound success rate of 37.18%, with Qwen3-Omni-30B achieving 19.71%. Figure 5(e) shows the progressive success rate. Interference from the concurrent sound event poses a major challenge to completion of this setting (see Section. 3.2).
@@
Qualitative Result. Figure 6 presents qualitative visualizations of the agent’s execution across the three proposed tasks. In each subfigure, the left panel depicts the top-down trajectory and sound source locations, while the right panel displays sequential third-person keyframes of the interaction. Figure 6 (a) shows the object relocation. Initially, the target object (an alarm) is located outside the agent’s field of view. Relying on sound signals for localization, the agent successfully navigates to the target and executes the Pick and Place skills. Figure 6 (b) illustrates the state transition. The agent accurately navigates to the sound-emitting door and performs the Open Door skill. Figure 6 (c) showcases the multi-event setting, which involves two simultaneous sound sources: a primary source (yellow) and a secondary source (green). Guided by the task planner, the agent prioritizes the primary source despite the acoustic interference. It first navigates to and opens the door. Once the primary source is interacted, the agent isolates the remaining secondary signal, navigates to the sink, and executes the Close Sink skill. More visualizations can be seen in supp.
Table 2. Effect of audio modality on navigate policy model. Text Cat. and Tgt. Pos. denote the text category and target position of the sound source, respectively.
Model Input Modality S.R. ($\uparrow$)
Audio Text Cat. Tgt. Pos.
I 0.00
II 0.00
III 57.66
IV 94.59
V 95.05
Table 3. Ablation Study on Input Modalities. We investigate the impact of input modalities and identify optimal combinations across task planning and three representative policy models. S.R. denotes success rate. (a) Task Planning: Investigates the impact of Head RGB and sound, finding that combining two modalities is the best. (b) Policy model for Navigate (Visual Modality): Using Head RGB and Head Depth yields the highest result. (c) Policy model for Pick and Open Door: Combining Head Depth and Arm Depth outperforms single-view inputs for both Pick and Open Door.
Method Input Modality S.R. ($\uparrow$)
Head RGB Sound
I 18.59
II 51.55
III 51.83
Model Input Modality S.R. ($\uparrow$)
Head RGB Head Depth
I 51.80
II 57.66
III 61.71
Model Input Modality S.R. ($\uparrow$)
Head Depth Arm Depth Pick Open Door
I 57.20 47.29
II 66.66 46.39
III 88.73 62.16
6.3. Ablation Study and Further Discussion
Effect of Audio Modality for Navigate Skill. Table 2 presents the impact of different input modalities for the navigate skill. Audio signals are converted into binaural spectrograms to provide implicit position information. Models lacking any spatial information (Model I and Model II) completely fail to navigate, yielding a 0.00% success rate. Relying solely on audio signals achieves a 57.66% success rate (Model III). Supplying the explicit target position directly reaches 94.59% (Model IV). Combining both audio and explicit target positions yields the best performance at 95.05% (Model V). Overall, these results indicate two key findings. First, spatial information of the sound source is crucial. Second, as an environmental signal, audio provides effective position information and is complementary to the explicit target position.
Effect of input Modality for task planning. Table 3(a) shows the influence of various input modalities for task planning in the multi-event setting. We utilize Qwen3-Omni-30B (Yang et al., 2025) as the baseline. Relying exclusively on Head RGB yields the lowest success rate at 18.59%. Utilizing only sound signals substantially elevates the performance to 51.55%. Integrating both Head RGB and sound attains the optimal result of 51.83%. Consequently, these findings show the importance of auditory information in task planning, while visual input from Head RGB serves a supplementary function.